$SPAD/src/input f02aaf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
ia:=4
n:=4
a:Matrix SF:=
 [[ 0.5 , 0.0 , 2.3 ,-2.6 ],_
  [ 0.0 , 0.5 ,-1.4 ,-0.7 ],_
  [ 2.3 ,-1.4 , 0.5 , 0.0 ],_
  [-2.6 ,-0.7 , 0.0 , 0.5 ]]
result:=f02aaf(ia,n,a,-1)

———-

References^{[1] nothing}

4